A ring and a disc are initially at rest, side by side, at the top of an inclined plane which makes
an angle 60° with the horizontal. They start to roll without slipping at the same instant of
time along the shortest path. If the time difference between their reaching the ground is
(2 − √3) /√10
1
Expert's answer
2019-08-05T15:19:29-0400

The acceleration of disc is given by formula


"a=\frac{g}{\sqrt {3}} (1)"

The acceleration of ring is given by formula


"a=\frac{g×\sqrt {3}}{4} (2)"

We can find time t

"t=\sqrt { \frac{2S}{a}} (3)"

where

"S=\frac{h} {\sin 60°} (4)"

The difference of time is equal to


"Δt=\frac{2 - \sqrt {3}} {\sqrt {10}} (5)"

We put (1) and (2) in (3) and we get using (5):

h=0.75 m


Answer

0.75 m